OpenStudy (learningallday):

What is the value of the expression? 19.2 + (–7 + 3.8) A. 16 B. 24 C. 29 D. 30

OpenStudy (learningallday):

@learningallday

OpenStudy (ahvei):

do you mind if i help you

OpenStudy (learningallday):

no i dont mind

OpenStudy (ahvei):

ok would you like me to explain or just give you the answer

OpenStudy (learningallday):

ether one it be find

OpenStudy (ahvei):

well D is the answer because if you add them up that is what you will get but if you would like i well walk you through the steps

OpenStudy (learningallday):

thank you so much

OpenStudy (ahvei):

you are very welcome is there any more questions

OpenStudy (learningallday):

yes

OpenStudy (learningallday):

What value of m makes the statement true? –(m + 15) = –18 + (–15) A. –33 B. –18 C. 3 D. 18

OpenStudy (ahvei):

B is the answer

OpenStudy (learningallday):

thank you

OpenStudy (ahvei):

tag me if you have anymore questions

OpenStudy (ahvei):

and your welcome

OpenStudy (learningallday):

What value of k makes the equation true? 12 + (–17) = –17 + k k =

OpenStudy (ahvei):

12 is the answer

OpenStudy (learningallday):

thats all thank you so much

OpenStudy (ahvei):

you are very welcome

OpenStudy (youngstudier):

It goes A. 16, D. 18, and the last one is 12. Most of the answers they gave you were wrong.

OpenStudy (ahvei):

plsss prove me wrong

OpenStudy (ahvei):

so explain it

OpenStudy (learningallday):

oh 1 16

OpenStudy (youngstudier):

-7 + 3.8 = -3.2 + 19.2 = 16

OpenStudy (ahvei):

ummm wrong because i did it by hand and i used the calculator

OpenStudy (learningallday):

i have more ?

OpenStudy (youngstudier):

–(m + 15) = –18 + (–15) It would be positive 18 because there is already a negative symbol outside the parentheses. -(18 + 15) = -33

OpenStudy (youngstudier):

I used a calculator, and I do 8th grade math, I'm pretty sure I know what I'm doing.

OpenStudy (learningallday):

(1 pt) What property is shown by the equation? Column A Column B 1. (34 + 11) + 9 = (11 + 34) + 9 2. 25 + (–15 + 4) = (25 + (–15)) + 4 3. –(–21 + 8) = 21 + (–8) 4. –30 + (–16) = –16 + (–30) A. commutative property B. associative property C. opposite of a sum property

OpenStudy (ahvei):

im in 12th grade dumby

OpenStudy (ahvei):

did you use calculator tab like i did

OpenStudy (youngstudier):

Really? Wow. No comment. Anything I'd say probably wouldn't be nice.

OpenStudy (ahvei):

sounds like me

OpenStudy (learningallday):

hello

OpenStudy (ahvei):

cause my mouth aint put on backwards i dont hold back no punches so ima warn u now TRY ME!!!!!!!!!!!!!!!

OpenStudy (learningallday):

can yall stop trash talking in help me please

OpenStudy (youngstudier):

Have @ahvei help you, because as a 12th grader, she so obviously, and I put stress of obviously, knows precisely what she is doing.

OpenStudy (youngstudier):

@learningallday , you entered their answers and got a 40? .....

OpenStudy (learningallday):

yup

OpenStudy (ahvei):

FOR 1 KEEP MY NAME OUT OF YOUR MOUTH FOR 2 NO ONE ASKED U TO COME HERE ANYWAYS UR retriceWAS FOLLOWING ME ANYWAYS BECAUSE U GOT MAD I HELPED THAT GIRL AND SHE PASSED

OpenStudy (youngstudier):

He/she failed their test.

OpenStudy (youngstudier):

80 and above is mastered, under 80 is failed.

OpenStudy (ahvei):

WHO BECUASE THAT OTHER PERSON I HELPED PASSED AND THEY HAD THE SAME QUESTIONS

OpenStudy (youngstudier):

No, they had different questions.

OpenStudy (ahvei):

AND THE PERSON I HELPED GOT A 100 AND WTF U TALKING ABOUT U DK THE PERSON

OpenStudy (ahvei):

YOU WASNT EVEN IN THE CHAT WITH ME

OpenStudy (youngstudier):

*Sigh* Oh wow! You're really gonna fight me over the internet? What's the worst you can do, caps-lock me to death?

OpenStudy (youngstudier):

@learningallday , open a new question and close this one.

OpenStudy (learningallday):

sure